a farmer with 4000 meters of fencing wants to encloce a rectangular plot that borders on a river.
*if the farmer does not fence the side along the river, what is the largest area that can be enclosed?
*provide the area function used and graph it. Identify the maximum and interpret what it means.

Answer by checkley77(12844)

THIS IS A SPECIAL CASE WHERE THE LENGTH IS TWICE THE WIDTH & THERE IS ONLY 1 LENGTH BUT 2 WIDTHS.
L=2W
2W+L=4,000
2W+2W=4,000
4W=4,000
W=4,000/4
W=1,000 ANS. FOR THE EACH OF THE WIDTHS.
L=2*1,000=2,000 ANS. FOR THE 1 LENGTH.
PROOF:
2*1,000+2,000=4,000
2,000+2,000=4,000
4,000=4,000
PROOF:
1,000*2,000=2,000,000 TOTAL AREA.
TEST:
ADD 2 TO THE LENGTH & SUBTRACT 1 FROM EACH OF THE SIDES.
2002*999=1.999,998
ADD 1 TO EACH OF THE WIDTHS & SUBTRACT 2 FROM THE LENGTH.
1998+1001=1,999,998